OpenStudy (anonymous):

Please help for medal? Marco and his two younger sisters would like to purchase a silver charm bracelet for their mother’s birthday. They went to the mall and found what they were looking for at Store A. In Store A, the bracelet without charms costs $85 and each charm costs $15 If Marco and his sisters have saved $250, make a graph to show all of the charms they could purchase with a bracelet if they had saved between $85 and $250.

OpenStudy (anonymous):

im here :) ok one minute

OpenStudy (anonymous):

Did you know "okay" is one of the most miss spelled english name out there? :O

OpenStudy (anonymous):

do you have geogebra?

OpenStudy (anonymous):

Oh god, is that a new math subject?

OpenStudy (anonymous):

NO lol

OpenStudy (anonymous):

its a software that flvs reccomends

OpenStudy (anonymous):

I'm already scared of geometry, and algebra was difficult enough for me. I don't want both subjects in one!

OpenStudy (anonymous):

Oh, thank god!

OpenStudy (anonymous):

go to the wedsite and download it can you can make graphs on it

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

website * i am misspelling everything today lol

OpenStudy (anonymous):

http://www.geogebra.org/cms/en/

OpenStudy (anonymous):

I'm sorry? Whats a "wedsite" is that a place where you can get married online?

OpenStudy (anonymous):

Yes, I created it and it has 125 people on it. :) it was created 6 years ago and we already have twelve people, how exciting

OpenStudy (anonymous):

What about this site?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

click on the link, it will take you to geogebra

OpenStudy (anonymous):

wedsite actually exists...

OpenStudy (anonymous):

Already done that

OpenStudy (anonymous):

rofl no it doesnt :O

OpenStudy (anonymous):

OH MY GOD IT DOES

OpenStudy (anonymous):

LETS JOIN IT! rofl

OpenStudy (anonymous):

LETS DO IT

OpenStudy (anonymous):

do wha?

OpenStudy (anonymous):

JOIN IT

OpenStudy (anonymous):

oh, But i'm already married :O

OpenStudy (anonymous):

:O

OpenStudy (anonymous):

to Julian Pizzahut Mcdonald :P

OpenStudy (anonymous):

Not too sound rude or anything but I need to finish 3 assignments by 7:30 and I only did 1. So I need to finish this and another exam. So if we could finish this quickly I'll give you 3 medals :O!!!!!!

OpenStudy (anonymous):

I'm leaving to go to Atlanta tonight

OpenStudy (anonymous):

AHA omg sry i forgot we were doings that

OpenStudy (anonymous):

Sorry if I'm sounding rude

OpenStudy (anonymous):

Is this site necessary?

OpenStudy (anonymous):

the graph, yes! okay um.. well i couldnt graph it so heres the descvription of it,

OpenStudy (anonymous):

um, well it works well for me..

OpenStudy (anonymous):

origin y-axis: 85 origin x-axis: 0 y-axis interval: 5 x-axis interval: 1

OpenStudy (anonymous):

thats the graph.. geogebra can be complicated sometimes so i couldnt make a graph like that... I am still learning how to use it

OpenStudy (anonymous):

So how do I graph this?

OpenStudy (anonymous):

f(x)=85+15(5) f(x)=85+75 f(x)=160

OpenStudy (anonymous):

do you have it open?

OpenStudy (anonymous):

Geogebra? Yes.

OpenStudy (anonymous):

Where do I find this?

OpenStudy (anonymous):

Do I need to download anything?

OpenStudy (anonymous):

yes, you need to downlaod geogebra

OpenStudy (anonymous):

downloading

OpenStudy (anonymous):

oh ok, well tell me when it is done downloading and ill tell you what to do

OpenStudy (anonymous):

for now lets do the other questions

OpenStudy (anonymous):

It finished instantly

OpenStudy (anonymous):

oh ok, open it up :)

OpenStudy (anonymous):

Instantly opened

OpenStudy (anonymous):

:O

OpenStudy (anonymous):

ill tell you what to click

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

does it show the grid?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok now, right click the mouse

OpenStudy (anonymous):

at the bottom of that gray pop up box, it should say graphics

OpenStudy (anonymous):

click on graphics and tell me what you see

OpenStudy (anonymous):

I am right clicking the grid but nothing is happening

OpenStudy (anonymous):

nothing it popping up?

OpenStudy (anonymous):

theres another way.

OpenStudy (anonymous):

@ImaGuest r u there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, look at the file i attached,

OpenStudy (anonymous):

go to the top right corner, you should see something that has arrows, and something like a gear?

OpenStudy (anonymous):

click on the gear. when you click on it, it should say graphics on one of those options

OpenStudy (anonymous):

yeah thats not even there

OpenStudy (anonymous):

Do I need to sign in?

OpenStudy (anonymous):

:O

OpenStudy (anonymous):

no you dont need to sign in

OpenStudy (anonymous):

look at the file, do you see it now?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

do you have windows 7?

OpenStudy (anonymous):

or just windows?

OpenStudy (anonymous):

Windows 7

OpenStudy (anonymous):

do you have something called snipping tool?

OpenStudy (anonymous):

its a program that comes with windows 7

OpenStudy (anonymous):

This is all that shows

OpenStudy (anonymous):

thats weird cuz mine looks way different and easier than urs... i think urs might be an updated version

OpenStudy (anonymous):

unfortunately since urs is different, i cant help with the graph. but i didnt make the graph for mine either so what i did was type a description of what the graph should look like

OpenStudy (anonymous):

I need to go in 7 minutes :O

OpenStudy (anonymous):

If Marco and his sisters have saved $250, make a graph to show all of the charms they could purchase with a bracelet if they had saved between $85 and $250. I had trouble making the graph, the program is sort of complicated so here is a description of what it should look like. origin y-axis 85 origin x-axis 0 y-axis interval 5 x-axis interval 1

OpenStudy (anonymous):

thats the answer i put in

OpenStudy (anonymous):

Marco found five charms that he likes. Using your function, determine the cost of the bracelet he would make. f(x)=85+15(5) f(x)=85+75 f(x)=160

OpenStudy (anonymous):

these are the answers

OpenStudy (anonymous):

Thank you, can you quickly help with D and E?

OpenStudy (anonymous):

Oh thats for D, then C and E?

OpenStudy (anonymous):

Marco is wondering from which store he should buy the items so that he spends the least amount of money. Which store should Marco use if he wanted the five charms and silver bracelet all together? What advice can you tell Marco to help him shop wisely depending on the number of charms he wants to buy? Justify how you reasoned your answers. A B f(x)=85+15(5) f(x)=100+12(5) f(x)=85+75 f(x)=100+60 f(x)=160 f(x)=160

OpenStudy (anonymous):

omg I didn't even scroll down till now. There is a whole other question ;(

OpenStudy (anonymous):

Ikr?

OpenStudy (anonymous):

its fine, ill get you through it